At a suburban shopping mall 70% of customers pay by ecocash.Find the probability that in a randomly selected sample of ten customers : exactly 3 pay by ecocash and more than 6 pay by ecocash

1
Expert's answer
2022-05-09T18:23:50-0400

Let "X=" the number of customers pay by ecocash:"X\sim Bin(n, p)."

Given "n=10, p=0.7, q=1-p=0.3."

1.


"P(X=3)=\dbinom{10}{3}(0.7)^{3}(0.3)^{10-3}=0.009001692"

2.


"P(X>6)=1-P(X=7)-P(X=8)"

"-P(X=9)-P(X=10)"

"=1-\dbinom{15}{7}(0.7)^{7}(0.3)^{10-7}-\dbinom{15}{8}(0.7)^{8}(0.3)^{10-8}"

"-\dbinom{15}{9}(0.7)^{9}(0.3)^{10-9}\dbinom{15}{10}(0.7)^{10}(0.3)^{10-10}"

"=0.6496107184"